# Self-Managing DRAM: A Low-Cost Framework for Enabling Autonomous and Efficient in-DRAM Operations

### v6

@@ 7.3 SMD’s Energy Consumption
Fig. 11 shows the average DRAM energy consumption normalized to the DDR4 baseline for single- and four-core workloads. We make two major observations.
@@
First, SMD-Combined and SMD-FR reduce DRAM energy by 3.9% (4.3%) and 5.2% (5.2%) across all single-core (4c-high) workloads compared to the baseline, respectively. SMD-Combined performs close to the hypothetical No-Refresh configuration, providing 59.7% of the energy reduction benefits of No-Refresh on average across 4c-high workloads. We attribute the reduction in DRAM energy to i) the reduced DRAM background energy consumption because SMD shortens the execution time for these workloads and ii) eliminating DRAM commands for maintenance operations (e.g., for periodic refresh) on the power-hungry DDRx bus. Second, SMD-FR-1LR increases average single-core (4c-high) workload DRAM energy consumption by 0.1% (3.0%). While SMD-FR-1LR also eliminates DRAM commands for maintenance operations, it induces energy overheads mainly because of the increased execution time for the evaluated workloads.
7.4 Hardware Overhead
An SMD chip needs to transmit ACT_NACK signals to the memory controller (MC). SMD can 1) repurpose the existing alert_n pin in DDR4 chips [78, 124] or 2) introduce an extra physical pin to transmit ACT_NACK signals.
1) alert_n. alert_n is currently used to inform the MC that the DRAM chip detected a CRC or parity check failure on the issued command, and thus the MC must issue the command again. The alert_n signal can simply be asserted not only on CRC or parity check failure, but also when the MC attempts accessing a row in a locked region. This approach does not require an additional physical pin. However, 1) assigning multiple meanings to an already ambiguous alert_n signal [78] could complicate MC design and 2) alert_n, as currently defined in the standard [78], is an open drain signal that is slow and not precise enough.
2) Introducing a new pin. To potentially simplify MC design, SMD can introduce a new pin to transmit ACT_NACK signals. Only one new pin between the MC and all memory devices (e.g., all DRAM chips in a memory channel or a rank) is enough for systems that use rank-based DRAM chip organizations (e.g., DDR4 DIMMs). For example, 128 new pins are needed for a high-end system with 32 memory channels equipped with 4-rank memory chips in each channel. However, a system of this scale already has 4’094 processor pins [125], and 128 new pins would amount to a relatively small 3.13% increase in pins. Similar to how alert_n signal works, the module can send the MC a single ACT_NACK when any of the per-chip ACT_NACK signals are asserted.
We individually discuss the other changes required on the existing DRAM chip and the MC circuitry.
DRAM Chip Modifications. We use CACTI [126] to evaluate the hardware overhead of the changes that SMD introduces over a conventional DRAM bank (highlighted in Fig. 2) assuming technology.121212Capitalizing on the latest DRAM technology libraries to implement the changes that SMD introduces would likely provide more accurate area overhead estimations. However, such libraries and tools are proprietary and we do not have access to them. Lock Region Table (LRT) is a small table that stores a single bit for each lock region in a DRAM bank to indicate whether or not the lock region is under maintenance. In our evaluation, we assume that a bank is divided into lock regions. Therefore, LRT consists of only bits, which are indexed using a 4-bit lock region address. According to our evaluation, an LRT incurs only area overhead per bank. The area overhead of all LRTs in a DRAM chip is only 0.001% of a DRAM chip. The access time of an LRT is , which is only 0.4% of typical row activation latency (tRCD) of .
SMD adds a per-region RA-latch to enable accessing one lock region while another is under maintenance. An RA-latch stores a pre-decoded row address, which is provided by the global row address decoder, and drives the local row decoders where the row address is fully decoded. This design builds on the basic design proposed in SALP [79] and refresh-access parallelization introduced in [16, 80]. According to our evaluation, all RA-latches incur a total area overhead of 1.6% of a DRAM chip. An RA-latch has only latency, which is negligible compared to tRCD.
Besides these changes that are the core of the SMD substrate, a particular maintenance mechanism may incur additional area overhead. We evaluate the DRAM chip area overhead of the maintenance mechanisms presented in section 5. The simple refresh mechanism, SMD-FR, requires only additional area in a DRAM chip. The DRAM chip area overhead of the refresh mechanism is less than 0.1% of a typical DRAM chip. SMD-DRP requires a large Counter Table with 1224 counters per bank for the RowHammer threshold value that we use in our performance evaluation. Across a DRAM chip, SMD-DRP requires area, which is 7.0% of a typical DRAM chip size. The control logic of SMD-MS is similar to the control logic of SMD-FR and it requires only additional area excluding the area of the ECC engine, which is already implemented by DRAM chips that support in-DRAM ECC.
Memory Controller Modifications. We slightly modify the MC’s scheduling mechanism to retry a rejected ACT command as we explain in section 4.3. Upon receiving an ACT_NACK, the MC marks the bank as precharged. An existing MC already implements control circuitry to pick an appropriate request from the request queue and issue the necessary DRAM command based on the DRAM bank state (e.g., ACT to a precharged bank or RD if the corresponding row is already open) by respecting the DRAM timing parameters. The ACT Retry Interval (ARI) is simply a new timing parameter that specifies the minimum time interval for issuing an ACT to a lock region after receiving an ACT_NACK from the same region. Therefore, SMD can be implemented in existing MCs with only slight modifications by leveraging the existing request scheduler (e.g., FRFCFS [115]).
Tracking locked regions. To apply the ARI timing parameter (e.g., to an ACT command targeting a locked region), the memory controller tracks which regions are locked. The memory controller could store the address of the locked region in every DRAM bank (since only one region can be under maintenance at a time in a bank). 1 rank address bit, 4 bank address bits, and 4 lock region address bits (9 bits) are sufficient to track every under-maintenance lock region in 2 ranks, 8 chips, and 16 banks (256 unique banks). Therefore, the storage cost for tracking locked regions is only 288 bytes for the evaluated memory channel with a dual-rank x8 memory module.
Address mapping schemes. SMD does not require modifications to how a physical address is mapped to a memory-controller-visible DRAM address. In section 7, we evaluate SMD using a mapping scheme that interleaves consecutive cache blocks across DRAM channels (from the most significant to the least significant bit, the physical address is mapped to rows, banks, ranks, columns, and channels). We also evaluate SMD using the mapping scheme described in [127] that aims to exploit bank-level parallelism. For this mapping scheme, SMD-FR provides 8.7% speedup over the baseline system, on average across 4c-high workloads. We leave detailed evaluation of SMD with different address mapping schemes for future work.
No further changes are required in the MC to support different maintenance mechanisms enabled by SMD. Thus, SMD enables DRAM designers to update existing maintenance mechanisms and implement new ones without any further changes to MCs that support SMD.
7.5 Sensitivity Studies
We analyze the performance and DRAM energy effects of different SMD configuration parameters.
7.5.1 DRAM Refresh Period
Fig. 12 plots the speedup and energy reduction that SMD-FR, SMD-VR, and SMD-combined achieve over DDR4 for different refresh periods across single- and four-core workloads. We make three key observations. First, the performance and energy benefits of SMD-FR/SMD-VR increase as the refresh period reduces, achieving 50.5%/53.6% speedup and 25.6%/29.0% DRAM energy reduction on average at refresh period across 4c-high workloads. Thus, SMD-based refresh mechanisms will become even more beneficial to employ in future DRAM chips that are expected to require more frequent refresh [17, 8, 86]. Second, SMD-combined provides 51.5% speedup and 27.9% DRAM energy reduction on average across 4c-high at refresh period, showing that the overheads of SMD-based RowHammer protection and memory scrubbing mechanisms remain low even at high refresh rates. Third, all SMD-based maintenance mechanism eliminate most of the performance overhead of refresh across all refresh periods (e.g., SMD-combined reaches 95.6% of the hypothetical NoRefresh DRAM at ).
We conclude that SMD-based maintenance mechanisms improve DRAM reliability and security while greatly reducing the overhead of DDR4 refresh. Their performance benefits are expected to increase for future DRAM chips that are likely to require high refresh rates.
7.5.2 ACT_NACK Divergence Across Chips
In Section 4.4, we explain divergence in SMD maintenance operations can happen when different DRAM chips in the same rank perform maintenance operation at different times. Such a divergence leads to a partial row activation when the activated row is in a locked region in some DRAM chips but not in others. To handle partial row activations, we develop three policies.
Precharge. With the Precharge policy, the MC issues a PRE command to close the partially activated row when some DRAM chips send ACT_NACK but others do not. After closing the partially activated row, the MC can attempt to activate a row from a different lock region in the same bank.
Wait. With the Wait policy, the MC issues multiple ACT commands until a partially activated row becomes fully activated. When some chips send ACT_NACK for a particular ACT but others do not, the MC waits for ARI and issues a new ACT to attempt activating the same row in DRAM chips that previously sent ACT_NACK.
Hybrid. We also design a Hybrid policy, where the memory controller uses the Precharge policy to close a partially activated row if the request queue contains or more requests that need to access rows in different lock regions in the same bank. If the requests queue has less than requests to different lock regions, the MC uses the Wait policy to retry activating the rest of the partially activated row.
In Fig. 13, we compare the performance and energy savings of SMD-FR and SMD-VR when using the three ACT_NACK divergence handling policies across 4c-high workloads. The plots show results for the common-case (CC) and worst-case (WC) scenarios with regard to when maintenance operations happen across different SMD chips in the same rank. In the common-case scenario, the DRAM chips generally refresh the same row at the same time due to sharing the same DRAM architecture design. However, refresh operations in some of the DRAM chips may still diverge during operation depending on the refresh mechanism that is in use. For example, SMD-VR refreshes retention-weak rows, whose locations may differ across the DRAM chips, at a higher rate compared to other rows, resulting in divergence in refresh operations across the DRAM chips in a rank. In the worst-case scenario, we deliberately configure the DRAM chips to refresh different rows at different times. For this, we 1) delay the first refresh operation in by , where and is the latency of a single refresh operation, and 2) set the Lock Region Counter (LRC) of to .
We make three key observations. First, the performance and DRAM energy consumption only slightly varies across different divergence handling policies in both common-case and worst-case scenarios. This is because, after a partial row activation happens, a different row that is not in a locked region in all off of the DRAM chips often does not exist in the memory request queue. As a result, the Precharge and Hybrid policies perform similarly to the Wait policy.
Second, SMD-FR performs worse than the DDR4 baseline in the worst-case refresh distribution scenario (i.e., average slowdown of 2.6% with the Wait policy). Certain individual workloads experience even higher slowdown (up to 12.8% with the Wait policy). The reason is that, when different DRAM chips lock a region at different times, the total duration during which the lock region is unavailable becomes times the duration when all chips simultaneously refresh the same lock region. This significantly increases the performance overhead of refresh operations.
Third, SMD-VR does not suffer much from the divergence problem and outperforms the DDR4 baseline even in the worst-case scenario. This is because SMD-VR mitigates the DRAM refresh overhead by significantly reducing the number of total refresh operations by exploiting retention-time variation across DRAM rows. Thus, the benefits of eliminating many unnecessary refresh operations surpass the overhead of ACT_NACK divergence.
We conclude that 1) although SMD-FR suffers from noticeable slowdown for the worst-case scenario, which should not occur in a well-designed system, it still provides comparable performance to conventional DDR4 and 2) SMD-VR outperforms the baseline and saves DRAM energy even in the worst-case scenario.
7.5.3 Comparison to Conventional Scrubbing
@@
Fig. 14 compares the performance and energy overheads of conventional DDR4 scrubbing to SMD-MS across 4c-high workloads. In the figure, DDR4 Scrubbing represents the performance and energy overhead of conventional scrubbing compared to DDR4 without memory scrubbing. Similarly, SMD-MS represents the performance and DRAM energy overhead of SMD-based scrubbing compared to SMD-FR.
@@
We make two observations. First, both DDR4 scrubbing and SMD-MS have negligible performance overhead for scrubbing periods of 5 minutes and larger because scrubbing operations are infrequent at such periods. Second, DDR4 scrubbing causes up to 1.49%/8.84% average slowdown for 1 minute/10 second scrubbing period (up to 2.37%/14.26%), while SMD-MS causes only up to 0.34%/1.78% slowdown. DDR4 scrubbing has high overhead at low scrubbing periods because moving data from DRAM to the MC to perform scrubbing is inefficient compared to performing scrubbing within DRAM using SMD-MS. Scrubbing at high rates may become necessary for future DRAM chips as their reliability characteristics continuously worsen. Additionally, mechanisms that improve DRAM performance at the cost of reduced reliability [15, 128] can use frequent DRAM scrubbing to achieve the desired DRAM reliability level.
We conclude that SMS performs memory scrubbing more efficiently than conventional MC based scrubbing and it enables scrubbing at high rates with small performance and energy overheads.
7.5.4 Comparison to PARA in Memory Controller
@@
7.5.5 SMD-DRP Maximum Activation Threshold
7.5.6 Number of Vulnerable Neighbor Rows
7.5.7 SMD’s Performance Sensitivity to Number of Cores
The four-core workloads with high memory intensity (4c-high) largely saturate the DRAM bandwidth offered by four channels. Therefore, four cores are sufficient to capture the performance impact of SMD for cases where DRAM bandwidth utilization is high. To demonstrate this, we evaluate the performance of SMD-FR for a refresh period using highly memory-intensive 8-core (i.e., 8c-high) workloads. The speedup is comparable to the speedup of 4c-high workloads. At 32ms refresh period, SMD-FR provides 8.1% average speedup for 8c-high workloads while it provides 8.7% speedup for 4c-high workloads. SMD provides a slightly lower speedup for 8c-high because of the increased memory request queueing delay, as eight-core workloads produce more memory requests than four-core workloads.
8 Discussion
8.1 Prioritizing Memory Requests
To minimize the impact of rejected activate commands on system performance, an SMD chip could apply a different activate command rejection policy that prioritizes activate commands over maintenance operations. We describe and estimate the performance impacts of one such policy.
“Pause Maintenenace” Policy (SMD-PMP). The key idea of SMD-PMP is to pause an ongoing maintenance operation for a lock region when the memory controller issues an activate command to the lock region. The SMD chip could resume the maintenance operation when the memory controller precharges the bank (i.e., finishes accessing the lock region). For example, the SMD chip sequentially refreshes eight rows in a lock region (section 5.1) during a periodic refresh maintenance operation. If the chip receives an activate command to this lock region while only four out of eight rows have been refreshed, the chip does not continue refreshing the fifth row, but yields control of the lock region to the memory controller.
Estimating the Performance of SMD-PMP. We model SMD-PMP-FR as SMD-FR that refreshes one DRAM row in a lock region before releasing the lock (i.e., SMD-PMP-FR locks a region for at most ). Fig. 18 shows weighted speedup for 60 four-core workloads (20 per memory intensity level) on the left and the number of ACT_NACKs over the number of issued activate commands (which we call “ACT_NACK rate”), averaged across all workloads in every intensity level on the right.
We make two key observations from Fig. 18. First, SMD-PMP-FR provides a larger average speedup than SMD-FR for 4c-high workloads. SMD-PMP-FR’s ACT_NACK rate is 2% smaller than SMD-FR’s, which explains the 0.4% larger average speedup it provides over SMD-FR. Second, SMD-FR-1LR induces a smaller average system slowdown than SMD-PMP-FR-1LR. We conclude that an SMD-PMP design (whose performance could be accurately estimated using our methodology) that pauses maintenance operations to serve activate commands can improve performance.

### v7

@@ 8.3 Energy Consumption
Fig. 10 shows the DRAM energy consumption (normalized to the DDR4 baseline) across all tested single-core and four-core workloads with SMD-FR, SMD-VR, SMD-FR + SMD-PRP, SMD-FR + SMD-DRP, SMD-FR + SMD-MS, SMD-Combined, and No-Refresh. The white circles (error lines) represent the average (minimum and maximum) normalized DRAM energy consumption. A lower value on the y-axis indicates smaller DRAM energy consumption. We make three major observations.
@@
First, all SMD configurations reduce DRAM energy consumption of all evaluated workloads compared to the baseline. We attribute the reduction in DRAM energy to i) the reduced DRAM background energy consumption because SMD shortens the execution time for these workloads and ii) the elimination of DRAM commands for maintenance operations (e.g., for periodic refresh) on the power-hungry DDRx bus. Second, SMD provides substantial DRAM energy savings for memory intensive workloads with relatively high LLC MPKI. For example, for 4c-medium and 4c-high workloads, on average, SMD-Combined reduces DRAM energy by 4.8% and 4.3%. For 4c-high workloads, SMD-Combined provides 59.6% of the average DRAM energy savings provided by No-Refresh, while also improving system robustness by conducting periodic refresh, RowHammer protection, and periodic scrubbing operations inside the DRAM chip. Third, SMD-VR’s energy savings (6.9% on average across all 4c-high workloads) are higher than other SMD configurations because SMD-VR mitigates DRAM refresh overhead.
We conclude that SMD, by autonomously performing maintenance operations inside the DRAM chip, significantly reduces DRAM energy consumption compared to the baseline system that issues DRAM commands over the power-hungry DDRx bus to perform maintenance operations.
8.4 Performance Comparison to DARP and DSARP
Fig. 9 shows weighted speedup (normalized to the weighted speedup of the DDR4 baseline) for 60 four-core workloads (20 per memory intensity level) with DARP-Combined, DSARP-Combined, and SMD-Combined. The white circles inside each box (error lines) represent the average (minimum and maximum) normalized weighted speedup across the 20 workloads in the corresponding group.
We observe that SMD-Combined outperforms both DARP- and DSARP-Combined on average across all workloads in each memory intensity category. For example, SMD-Combined provides 8.6% and 4.1% speedups on average across 4c-high workloads over DARP-Combined and DSARP-Combined, respectively. We attribute these speedups to i) SMD’s maintenance-access parallelization and ii) SMD’s ability to perform maintenance operations autonomously inside the DRAM chip without the memory controller having to issue DRAM commands. While DSARP-Combined can concurrently access main memory and perform a maintenance operation, it needs to issue a DRAM command (e.g., a per bank refresh command for periodic refresh) for each maintenance operation and delay other DRAM commands that the memory controller issues to serve main memory requests. In contrast, SMD-Combined autonomously performs each maintenance operation inside the DRAM chip and does not incur delays for DRAM commands (except those incurred by ACT_NACKs). DARP-Combined cannot concurrently perform a maintenance operation and access main memory, thereby DARP-Combined performs worse than both SMD-Combined and DSARP-Combined.
We conclude that SMD outperforms a system that allows the memory controller to concurrently perform a maintenance operation and a memory access.
9 Design Choices and Sensitivity Analyses
We 1) describe and evaluate an SMD policy that can prioritize activate commands targeting a locked region over maintenance operations, 2) analyze SMD’s performance and energy sensitivities to the number of lock regions, refresh period, number of cores in the system, memory scrubbing rate, RowHammer thresold, and RowHammer blast radius, 3) compare SMD-based probabilistic RowHammer protection’s (SMD-PRP) performance and energy to MC-based probabilistic RowHammer protection’s (PARA [12]).
9.1 Prioritizing Memory Requests
To minimize the impact of rejected activate commands on system performance, an SMD chip could apply a different activate command rejection policy that prioritizes activate commands over maintenance operations either selectively or globally. We describe and estimate the performance impact of one such policy.
“Pause Maintenance” Policy (SMD-PMP). The key idea of SMD-PMP is to pause an ongoing maintenance operation for a lock region when the memory controller issues an activate command to the locked region. The SMD chip could resume the maintenance operation when the memory controller precharges the bank (i.e., finishes accessing the lock region). For example, the SMD chip sequentially refreshes eight rows in a lock region (section 5.1) during a periodic refresh. If the chip receives an activate command to this lock region while only four out of eight rows have been refreshed, the chip does not continue refreshing the fifth row, but yields control of the lock region to the memory controller.
Estimating the Performance of SMD-PMP. We model SMD-PMP-FR (building on SMD-FR section 5.1) that refreshes only one DRAM row in a lock region before releasing the lock (i.e., SMD-PMP-FR locks a region for at most ). Fig. 12 shows normalized weighted speedup of 60 four-core workloads (20 per memory intensity level) with SMD-FR and SMD-PMP-FR on the left and the number of ACT_NACKs over the number of issued activate commands (which we call “ACT_NACK rate”), averaged across all workloads in every intensity level on the right. SMD-FR is configured as described in section 7 and it refreshes 8 DRAM rows in a lock region before releasing the lock.
We observe that SMD-PMP-FR provides a larger average speedup than SMD-FR for 4c-high workloads. SMD-PMP-FR’s ACT_NACK rate is 2% smaller than SMD-FR’s, which explains the 0.4% larger average speedup it provides over SMD-FR. We conclude that an SMD-PMP design that pauses maintenance operations to serve activate commands could improve performance.
9.2 Sensitivity to Number of Lock Regions
To understand SMD’s sensitivity to the number of lock regions in a DRAM bank, we evaluate the performance and DRAM energy of SMD-FR (that refreshes 8 DRAM rows in a lock region before releasing the lock) as we vary the number of lock regions in a bank. Fig. 13 (top) shows normalized weighted speedup across 20 four-core high memory intensity workloads (y-axis) for different number of lock regions (x-axis) with SMD-FR and Fig. 13 (bottom) shows normalized DRAM energy across the same workloads and the number of lock regions.
We make two key observations from Fig. 13. First, SMD-FR provides higher speedup from 1.8% to 10.0% (averaged across all 20 workloads) as the number of lock regions increases from 2 to 256. This is because more lock regions allow SMD to concurrently perform more maintenance operations and memory accesses in a bank. Increasing the number of lock regions beyond 16 provides diminishing returns as 16 lock regions per bank provides 88.8% of the speedup provided by 256 lock regions. Second, SMD-FR with only one lock region per bank causes 3.9% average slowdown across all workloads. SMD-FR with one lock region cannot concurrently perform maintenance operations and memory accesses in a bank because every maintenance operation locks the whole bank. We attribute the overheads of SMD-FR with one lock region per bank to the relatively high rate of ACT_NACK commands (not shown in the figure) it issues: This configuration of SMD issues an ACT_NACK command for every 11.7 activate commands, on average across all 4c-high workloads. SMD-FR with 16 lock regions per bank, in contrast, issues an ACT_NACK command only every 134.6 activate commands, on average across all 4c-high workloads. We make similar observations for DRAM energy.
We conclude that SMD with 16 lock regions in a bank yields robust performance and DRAM energy improvements at relatively low DRAM chip area cost (see section 6).
9.3 Sensitivity to DRAM Refresh Period
We evaluate SMD for refresh periods of , , , and . We observe that the performance benefit of SMD-FR increases as the refresh period reduces, achieving 3.6 speedup at refresh period averaged across 4c-high workloads. As the refresh period becomes smaller, the baseline system more frequently performs time-intensive periodic refresh operations (each of which takes =). At a refresh period of , the memory controller issues a command every of . Thus, in the baseline system, a DRAM rank is busy 70.9% of the time doing refresh. SMD also more frequently performs refresh operations (and more frequently ACT_NACK’s activate commands) as the refresh period reduces. However, SMD alleviates the performance overhead of refresh operations by periodically refreshing only one lock region at a time while allowing ACT commands targeting other lock regions to proceed in parallel.
We also evaluate a baseline system with per-bank refresh at a refresh period. Because a bank is busy 70.9% of the time doing refresh operations, SMD provides 1.7 average speedup across 4c-high workloads over the baseline system with per-bank refresh.
We conclude that SMD’s performance benefits will likely increase for future DRAM chips that require higher refresh rates as 1) the number of rows in a DRAM bank increases to improve DRAM density or 2) the refresh period reduces to improve error tolerance for shrinking DRAM technology or increasing operating temperatures with tighter DRAM-system integration (e.g., high bandwidth memory [142, 144]).
9.4 Sensitivity to Number of Cores
We evaluate the performance of SMD-FR for a refresh period using highly memory-intensive 1-, 2-, 4-, and 8-core workloads. At a 32ms refresh period, SMD-FR provides 8.0%, 8.0%, 8.7%, and 8.1% average normalized weighted speedup for 1-, 2-, 4-, and 8-core high memory intensity workloads, respectively. We conclude that SMD-FR provides substantial performance improvements across 1-, 2-, 4-, and 8-core workloads. We attribute the difference between the provided performance improvement across different numbers of cores to workload and memory access interleaving differences.
9.5 SMD- vs. Memory-Controller-Based Scrubbing
@@ 9.4 Sensitivity to Number of Cores
Fig. 14 compares the performance overheads of memory-controller-based memory scrubbing to SMD-based memory scrubbing (SMD-MS) across 4c-high workloads. We show performance overhead normalized to the baseline system.
@@
We make two observations. First, both MC-based scrubbing and SMD-MS have small performance overheads for scrubbing periods of 5 minutes and larger because scrubbing operations are infrequent at such periods. Second, MC-based scrubbing causes 1.5%/8.8% average slowdown for a 1 minute/10 second scrubbing period (up to 2.4%/14.3%), while SMD-MS causes only 0.2/1.1% average (up to 0.3%/1.8%) slowdown. MC-based scrubbing has high overhead at low scrubbing periods because moving data from DRAM to the MC to perform scrubbing is inefficient compared to performing scrubbing completely inside the DRAM chip using SMD-MS. Scrubbing at high rates may become necessary for future DRAM chips as their reliability characteristics continuously worsen. We conclude that SMD-MS performs memory scrubbing more efficiently than conventional MC-based scrubbing and it enables scrubbing at high rates with small performance overheads.
9.6 Sensitivity to RowHammer Threshold
9.7 Sensitivity to RowHammer Blast Radius
9.8 SMD-based Probabilistic
RowHammer Protection vs. PARA [12]